Calculate the area of triangle ABC with altitude, CD, given:
A(–6, –4), B(6, 5), C(–1, 6), and D(2, 2).
Round your answer to the nearest tenth if necessary.

Answer by Boreal(15235)

Base is distance between A and B (draw this)
that is sqrt (12^2+9^2)= sqrt(225) or 15
CD has distance sqrt (9+16) or sqrt(25)=5
Area is (1/2) bh=(1/2)*75=37.5 units^2